Identify the type of conic, if any, represented by this equation. If the equation does represent a conic, state its vertex or center.
5x^2+10x+8y^2-48y+37=0

Answer by stanbon(75887)

Identify the type of conic, if any, represented by this equation. If the equation does represent a conic, state its vertex or center.
5x^2+10x+8y^2-48y+37=0
Complete the square on the x-terms and on the y-terms.----
5(x^2+2x+1) + 8(y^2-6y+9) = -37+5*1+8*9
-----
5(x+1)^2 + 8(y-3)^2 = 40
----
(x+1)^2/8 + (y-3)^2/5 = 1
-----
Ellipse:
Center at (-1,3)
=========================
Cheers,
Stan H.
====================